tigerlover (tigerlover):

A 10 foot ladder is leaning against a building. The ladder makes a 45 degree angle with the building. How far up the building does the ladder reach? A. feet B. 5 feet C. feet D. feet

tigerlover (tigerlover):

\[A. 20\sqrt{2}\]

tigerlover (tigerlover):

\[C. 10\sqrt{2}\]

tigerlover (tigerlover):

\[D. 5\sqrt{2}\]

tigerlover (tigerlover):

@kgrendel0324

OpenStudy (kgrendel0324):

wait

tigerlover (tigerlover):

yes?